OpenStudy (anonymous):

Simplify the expression: 0.3x^3 + 0.6x^3 - 0.4x^3

OpenStudy (anonymous):

0.3+0.6 -0.4 = 0.5 so 0.5x^3 you can combine like terms because they are all x^3

OpenStudy (anonymous):

so I am basically adding 3 + 6 =9 - 4 = 5 and add the like terms of x^3 ? for 0.5x^3...??

OpenStudy (anonymous):

yeah basically thats it as long as its a single variable with the same exponent you can combine them

OpenStudy (anonymous):

I have 2 ore problems to ask assistance with. Is this where I type them in at?

OpenStudy (anonymous):

Use distributive property to write the expression without parentheses: 3(a+b)